Given a number 𝑑 = 𝑑𝑒𝑡(𝐴1, ⋯ , 𝐴𝑛 ), find 𝑑𝑒𝑡(𝐴𝑛, ⋯ , 𝐴1 ).


1
Expert's answer
2021-08-04T17:06:58-0400

Answer:-

Property of determinant



"det(A_1, ..., A_i, ..., A_j, ..., A_n)=-det(A_1, ..., A_j, ..., A_i ..., A_n)"

We swap "A_1" and "A_n," "A_2" and "A_{n-1},..."

If "n=4k-3" or "n=4k,k=1,2,...," then



"det(A_1, ..., A_n)=det(A_n,...A_1)"

If "n=4k-2" or "n=4k-1,k=1,2,...," then



"det(A_1, ..., A_n)=-det(A_n,...A_1)"